The focal length of a thin lens obtained by repeating measurements is given as follows: 10.1, 10.3, 9.9, 10.0, 9.8cm. what is the average focal length and the mean deviation.

1
Expert's answer
2021-11-16T09:54:30-0500

Average:


"a=\frac{10.1+10.3+9.9+10+9.8}{5}=10\text{ cm}."

Mean deviation:


"m=\frac{(10.1-a)+(10.3-a)+(a-9.9)+(10-a)+(a-9.8)}{5}=0.14\text{ cm}."